# Deepfake Generation and Detection: A Benchmark and Survey

### v3

@@ 3.1.4 Facial Attribute Editing
In this section, we review current methods chronologically, following the progression in overcoming technical challenges, primarily focusing on multiple attribute editing methods utilizing GANs. Finally, we summarize methods in Tab. V.
@@
$\bullet$ Comprehensive Editing. Facial attribute editing aims to selectively alter specific facial attributes without affecting others. Therefore, disentangling different facial attributes is a primary challenge. Early facial attribute editing models [289, 273] often achieve editing for a single attribute through data-driven training. For instance, Shen et al[289] propose learning the difference between pre-/post-operation images, represented as residual images, to achieve attribute-specific operations. However, single-attribute editing falls short of meeting expectations, and compression steps in the process often lead to a significant loss of image resolution, a common issue in early methods. The fundamental challenge in comprehensive editing and unrelated attribute modification is achieving complete attribute disentanglement. Many approaches [275, 276, 277, 290] have explored this. E.g., HifaFace [277] identifies cycle consistency issues as the cause of facial attribute information loss that proposes a wavelet-based method for high-fidelity face editing, while TransEditor [290] introduces a dual-space GAN structure based on the transformer framework that improves image quality and attribute editing flexibility.
@@
$\bullet$ Irrelevant-attribute Retained. Another critical aspect of face editing is retaining as much target image information as possible in the generated images [276, 291, 21]. GuidedStyle [291] leverages attention mechanisms in StyleGAN [5] for the adaptive selection of style modifications for different image layers. IA-FaceS [21] embeds the face image to be edited into two branches of the model, where one branch calculates high-dimensional component-invariant content embedding to capture facial details, and the other branch provides low-dimensional component-specific embedding for component operations. Additionally, some approaches [279, 29, 28, 282] combine GANs with NeRF [26] for enhanced spatial awareness capabilities. Specifically, FENeRF [279] uses two decoupled latent codes to generate corresponding facial semantics and textures in a 3D volume with spatial alignment sharing the same geometry. CIPS-3D++ [282] enhances the model’s training efficiency with a NeRF-based shallow 3D shape encoder and an MLP-based deep 2D image decoder.
@@
$\bullet$ Text Driven facial attribute editing is a crucial application scenario and a recent hot topic in academic research [20, 292, 293, 283]. TextFace [292] introduces text-to-style mapping, directly encoding text descriptions into the latent space of pre-trained StyleGAN. TG-3DFace [283] introduces two text-to-face cross-modal alignment techniques, including global contrastive learning and fine-grained alignment modules, to enhance the high semantic consistency between the generated 3D face and the input text.
@@
$\bullet$ Diffusion-based models have been introduced into facial attribute editing [278, 280, 285, 286] and achieve excellent results. Huang et al[280] propose a collaborative diffusion framework, utilizing multiple pre-trained unimodal diffusion models together for multimodal face generation and editing. DiffusionRig [285] conditions the initial 3D face model, which helps preserve facial identity information during personalized editing of facial appearance based on the general facial details prior to the dataset.
@@
TABLE VI: Overview of representative forgery detection methods. Notations: ➀ FF++, ➁ DFDC, ➂ Celeb-DF, ➃ Deeperforensics, ➄ Self-build, ➅ UADFV, ➆ Celeb-HQ, ➇ DFDCp, ➈ FFHQ, ➉ DFD. Full table can be viewed in Appendix B.
@@
Space Domain Face X-ray [294] CVPR’20 ➀➁➂➉ Focusing on boundary artifacts of face fusion for forgery detection.
@@
Nirkin et al[295] TPAMI’21 ➀➁➂ Detecting swapped faces by comparing the facial region with its context (non-facial area).
@@
SBIs [296] CVPR’22 ➁➂➇➉ The belief that the more difficult to detect forged faces typically contain more generalized traces of forg -ery can encourage the model to learn a feature representation with greater generalization ability.
@@
RECCE [297] CVPR’22 ➀➁➂[92] Reconstruction learning on real samples to learn common compressed representations of real images.
@@
LGrad [30] CVPR’23 The gradient is utilized to present generalized artifacts that are fed into the classifier to determine the truth of the image.
@@
NoiseDF [298] AAAI’23 ➀➁➂➃ Extracting noise traces and features from cropped faces and background squares in video frames.
@@
Ba et al[299] AAAI’24 ➀➁➂ ➀➁➂ Multiple non-overlapping local representations are extracted from the image for forgery detection. A local information loss function, based on information bottleneck theory, is proposed for constraint.
@@
Time Domain FTCN [300] ICCV’21 ➀➁➂➃[182] It is believed that most face video forgeries are generated frame by frame. As each altered face is inde -pendently generated, this inevitably leads to noticeable flickering and discontinuity.
@@
LipForensics [301] CVPR’21 ➀➁➂ Concern about temporal inconsistency of mouth movements in videos.
@@
M2TR [108] ICMR’22 ➀➁➂➉ Capturing local inconsistencies at different scales for forgery detection using a multiscale transformer.
@@
Gu et al[302] AAAI’22 ➀➁➂[92] By densely sampling adjacent frames to pay attention to the inter-frame image inconsistency.
@@
Yang et al[34] TIFS’23 ➀➁➂ ➀➁➂ Treating detection as a graph classification problem and focusing on the relationship between the local image features across different frames.
@@
AVoiD-DF [39] TIFS’23 ➁➄[84] ➁➄[84] Multimodal forgery detection using audiovisual inconsistency.
@@
Choi et al[303] CVPR’24 ➀➂➃ ➀➂ Focus on the inconsistency of the style latent vectors between frames.
@@
Xu et al[304] IJCV’24 ➀➁➂➃ ➀➁➂➃ Forgery detection is conducted by converting video clips into thumbnails containing both spatial and temporal information.
@@
Frequency Domain F3-Net [36] ECCV’20 A two-branch frequency perception framework with a cross-attention module is proposed.
@@
FDFL [37] CVPR’21 Propose an adaptive frequency feature generation module to extract differential features from different frequency bands in a learnable manner.
@@
HFI-Net [305] TIFS’22 ➁➂➃➅[77] Notice that the forgery flaws used to distinguish between real and fake faces are concentrated in the mid- and high-frequency spectrum.
@@
Guo et al[306] TIFS’23 ➀➁ ➀➁➂ Designing a backbone network for Deepfake detection with space-frequency interaction convolution.
@@
Tan et al[307] AAAI’24 ➀➄ A lightweight frequency-domain learning network is proposed to constrain classifier operation within the frequency domain.
@@
Data Driven Dang et al[308] CVPR’20 ➂➅ Utilizing attention mechanisms to handle the feature maps of the detection model.
@@
Zhao et al[309] ICCV’21 ➀➁➂➃➇➉ Proposes pairwise self-consistent learning for training CNN to extract these source features and detect deep vacation images.
@@
Finfer [310] AAAI’22 ➀➂➇[92] Based on an autoregressive model, using the facial representation of the current frame to predict the facial representation of future frames.
@@
Huang et al[109] CVPR’23 ➀➁➂➉[182] A new implicit identity-driven face exchange detection framework is proposed.
@@
HiFi-Net [311] CVPR’23 Converting forgery detection and localization into a hierarchical fine-grained classification problem.
@@ 3.2 Forgery Detection
In this section, we review current forgery detection techniques based on the type of detection cues, categorizing them into three: Space Domain (Sec. 3.2.1), Time Domain (Sec. 3.2.2), Frequency Domain (Sec. 3.2.3), and Data-Driven (Sec. 3.2.4). We also summarize the detailed information about popular methods in Tab. VI.
@@ 3.2.1 Space Domain
$\bullet$ Image-level Inconsistency. The generation process of forged images often involves partial alterations rather than global generation, leading to common local differences in non-globally generated forgery methods. Therefore, some methods focus on differences in image spatial details as criteria for determining whether an image is forged, such as color [31], saturation [312], artifacts [313, 296, 297], gradient variations [30], etc. Specifically, RECCE [297] considers shadow generation from a training perspective, utilizing the learned representations on actual samples to identify image reconstruction differences. LGrad [30] utilizes a pre-trained transformation model, converting images to gradients to visualize general artifacts and subsequently classifying based on these representations. In addition, some works focus on detection based on differences in facial and non-facial regions [295], as well as the fine-grained details of image textures [314, 315]. Recently, Ba et al[299] focus not only on the discordance in a single image region but also on the detection of fused local representation information from multiple non-overlapping areas.
$\bullet$ Local Noise Inconsistency. Image forgery may involve adding, modifying, or removing content in the image, potentially altering the noise distribution in the image. Detection methods based on noise aim to identify such local or even global differences in the image. Zhou et al[32] propose a dual-stream structure, combining GoogleNet with a triplet network to focus on tampering artifacts and local noise in images. Nguyen et al[316] utilize capsule networks to detect forged pictures and videos in various forgery scenarios. NoiseDF [298] specializes in identifying underlying noise traces left behind in Deepfake videos, introducing an efficient and novel Multi-Head Relative Interaction with depth-wise separable convolutions to enhance detection performance.
@@ 3.2.2 Time Domain
$\bullet$ Abnormal Physiological Information. Forgery videos often overlook the authentic physiological features of humans, failing to achieve overall consistency with authentic individuals. Therefore, some methods focus on assessing the plausibility of the physiological features of the generated faces in videos. Li et al[91] detect blinking and blink frequency in videos as criteria for determining the video’s authenticity. Yang et al[94] focuses on the inconsistency of head poses in videos, comparing the differences between head poses estimated using all facial landmarks and those estimated using only the landmarks in the central region.
@@
$\bullet$ Inter-Frame Inconsistency. Methods [300, 302, 33, 34, 303, 304] based on inter-frame inconsistency for forgery detection aim to uncover differences in images between adjacent frames or frames with specific temporal spans. Gu et al[302] focus on inter-frame image inconsistency by densely sampling adjacent frames, while Yin et al[33] design a Dynamic Fine-grained Difference Capturing module and a Multi-Scale Spatio-Temporal Aggregation module to cooperatively model spatio-temporal inconsistencies. Yang et al[34] approach DeepFake detection as a graph classification problem, emphasizing the relationship information between facial regions to capture the relationships among local features across different frames. Choi et al[303] discover that the style variables in each frame of Deepfake work change. Based on this, they developed a style attention module to focus on the inconsistency of the style latent variables between frames.
@@
$\bullet$ Multimodal Inconsistency. The core idea behind multimodal detection algorithms is to make judgments based on the flow of prior information from multiple attributes rather than solely considering the image or audio differences of individual characteristics in each frame. The consideration of audio-visual modal inconsistency has received extensive research in various methods [38, 317, 318, 39]. POI-Forensics [317] proposes a deep forgery detection method based on audio-visual authentication, utilizing contrastive learning to learn the most distinctive embeddings for each identity in moving facial and audio segments. AVoiD-DF [39] embeds spatiotemporal information in a spatiotemporal encoder and employs a multimodal joint decoder to fuse multimodal features and learn their inherent relationships. Subsequently, a cross-modal classifier is applied to detect disharmonious operations within and between modalities. Agarwal et al[319] describe a forensic technique for detecting fake faces using static and dynamic auditory ear characteristics. Indeed, multimodal detection methods are currently a hotspot in forgery detection research.

### v4

@@ 3.1.4 Facial Attribute Editing
In this section, we review current methods chronologically, following the progression in overcoming technical challenges, primarily focusing on multiple attribute editing methods utilizing GANs. Finally, we summarize methods in Table 5.
@@
$\bullet$ Comprehensive Editing. Facial attribute editing aims to selectively alter specific facial attributes without affecting others. Therefore, disentangling different facial attributes is a primary challenge. Early facial attribute editing models shen2017learning ; zhou2017genegan often achieve editing for a single attribute through data-driven training. For instance, Shen et alshen2017learning propose learning the difference between pre-/post-operation images, represented as residual images, to achieve attribute-specific operations. However, single-attribute editing falls short of meeting expectations, and compression steps in the process often lead to a significant loss of image resolution, a common issue in early methods. The fundamental challenge in comprehensive editing and unrelated attribute modification is achieving complete attribute disentanglement. Many approaches shen2020interpreting ; yao2021latent ; gao2021high ; xu2022transeditor have explored this. E.g., HifaFace gao2021high identifies cycle consistency issues as the cause of facial attribute information loss that proposes a wavelet-based method for high-fidelity face editing, while TransEditor xu2022transeditor introduces a dual-space GAN structure based on the transformer framework that improves image quality and attribute editing flexibility.
@@
$\bullet$ Irrelevant-attribute Retained. Another critical aspect of face editing is retaining as much target image information as possible in the generated images yao2021latent ; hou2022guidedstyle ; huang2023ia . GuidedStyle hou2022guidedstyle leverages attention mechanisms in StyleGAN karras2019style for the adaptive selection of style modifications for different image layers. IA-FaceS huang2023ia embeds the face image to be edited into two branches of the model, where one branch calculates high-dimensional component-invariant content embedding to capture facial details, and the other branch provides low-dimensional component-specific embedding for component operations. Additionally, some approaches sun2022fenerf ; zhang2022fdnerf ; jiang2022nerffaceediting ; zhou2023cips combine GANs with NeRF mildenhall2020nerf for enhanced spatial awareness capabilities. Specifically, FENeRF sun2022fenerf uses two decoupled latent codes to generate corresponding facial semantics and textures in a 3D volume with spatial alignment sharing the same geometry. CIPS-3D++ zhou2023cips enhances the model’s training efficiency with a NeRF-based shallow 3D shape encoder and an MLP-based deep 2D image decoder.
@@
$\bullet$ Text Driven facial attribute editing is a crucial application scenario and a recent hot topic in academic research sun2022anyface ; hou2022textface ; aneja2023clipface ; yu2023towards . TextFace hou2022textface introduces text-to-style mapping, directly encoding text descriptions into the latent space of pre-trained StyleGAN. TG-3DFace yu2023towards introduces two text-to-face cross-modal alignment techniques, including global contrastive learning and fine-grained alignment modules, to enhance the high semantic consistency between the generated 3D face and the input text.
@@
$\bullet$ Diffusion-based models have been introduced into facial attribute editing preechakul2022diffusion ; huang2023collaborative ; ding2023diffusionrig ; kim2023diffusion and achieve excellent results. Huang et alhuang2023collaborative propose a collaborative diffusion framework, utilizing multiple pre-trained unimodal diffusion models together for multimodal face generation and editing. DiffusionRig ding2023diffusionrig conditions the initial 3D face model, which helps preserve facial identity information during personalized editing of facial appearance based on the general facial details before the dataset.
@@
Table 6: Overview of representative forgery detection methods. Notations: ➀ FF++, ➁ DFDC, ➂ Celeb-DF, ➃ Deeperforensics, ➄ Self-build, ➅ UADFV, ➆ Celeb-HQ, ➇ DFDCp, ➈ FFHQ, ➉ DFD.
@@
Space Domain Gram-Net liu2020global CVPR’20 ➆➈ ➆➈ The method posits that genuine faces and fake faces exhibit inconsistencies in texture details.
@@
Face X-ray li2020face CVPR’20 ➀➁➂➉ Focusing on boundary artifacts of face fusion for forgery detection.
Zhao et alzhao2021multi CVPR’21 ➀➁➂ A texture enhancement module, an attention generation module, and a bilinear attention pooling mod -ule are proposed to focus on texture details.
@@
Nirkin et alnirkin2021deepfake TPAMI’21 ➀➁➂ Detecting swapped faces by comparing the facial region with its context (non-facial area).
@@
SBIs shiohara2022detecting CVPR’22 ➁➂➇➉ The belief that the more difficult to detect forged faces typically contain more generalized traces of forg -ery can encourage the model to learn a feature representation with greater generalization ability.
@@
RECCE cao2022end CVPR’22 ➀➁➂zi2020wilddeepfake Reconstruction learning on real samples to learn common compressed representations of real images.
@@
LGrad tan2023learning CVPR’23 The gradient is utilized to present generalized artifacts that are fed into the classifier to determine the truth of the image.
@@
NoiseDF wang2023noise AAAI’23 ➀➁➂➃ Extracting noise traces and features from cropped faces and background squares in video frames.
@@
Ba et alba2024exposing AAAI’24 ➀➁➂ ➀➁➂ Multiple non-overlapping local representations are extracted from the image for forgery detection. A local information loss function, based on information bottleneck theory, is proposed for constraint.
Time Domain Yang et alyang2019exposing ICASSP’19 Focusing on the inconsistency in the head pose in videos by comparing the estimated head pose using all facial landmarks with the one estimated using only the landmarks in the central region.
@@
FTCN zheng2021exploring ICCV’21 ➀➁➂➃li2019faceshifter It is believed that most face video forgeries are generated frame by frame. As each altered face is inde -pendently generated, this inevitably leads to noticeable flickering and discontinuity.
@@
LipForensics haliassos2021lips CVPR’21 ➀➁➂ Concern about temporal inconsistency of mouth movements in videos.
@@
M2TR wang2022m2tr ICMR’22 ➀➁➂➉ Capturing local inconsistencies at different scales for forgery detection using a multiscale transformer.
@@
Gu et algu2022delving AAAI’22 ➀➁➂zi2020wilddeepfake By densely sampling adjacent frames to pay attention to the inter-frame image inconsistency.
@@
Yang et alyang2023masked TIFS’23 ➀➁➂ ➀➁➂ Treating detection as a graph classification problem and focusing on the relationship between the local image features across different frames.
@@
AVoiD-DF yang2023avoid TIFS’23 ➁➄khalid2021fakeavceleb ➁➄khalid2021fakeavceleb Multimodal forgery detection using audiovisual inconsistency.
@@
Choi et alchoi2024exploiting CVPR’24 ➀➂➃ ➀➂ Focus on the inconsistency of the style latent vectors between frames.
@@
Xu et alxu2024towards IJCV’24 ➀➁➂➃ ➀➁➂➃ Forgery detection is conducted by converting video clips into thumbnails containing both spatial and temporal information.
Peng et alpeng2024deepfakes TIFS’24 ➀➂➇ ➀➂➇ Focuse on inter-frame gaze angles, extracting gaze informations and employing spatio-temporal feature aggregation to combine temporal, spatial, and texture features for detection and classification.
@@
Frequency F3-Net qian2020thinking ECCV’20 A two-branch frequency perception framework with a cross-attention module is proposed.
@@
FDFL li2021frequency CVPR’21 Propose an adaptive frequency feature generation module to extract differential features from different frequency bands in a learnable manner.
@@
HFI-Net miao2022hierarchical TIFS’22 ➁➂➃➅korshunov2018deepfakes Notice that the forgery flaws used to distinguish between real and fake faces are concentrated in the mid- and high-frequency spectrum.
@@
Guo et alguo2023constructing TIFS’23 ➀➁ ➀➁➂ Designing a backbone network for Deepfake detection with space-frequency interaction convolution.
@@
Tan et altan2024frequency AAAI’24 ➀➄ A lightweight frequency-domain learning network is proposed to constrain classifier operation within the frequency domain.
@@
Data Driven Dang et aldang2020detection CVPR’20 ➂➅ Utilizing attention mechanisms to handle the feature maps of the detection model.
@@
Zhao et alzhao2021learning ICCV’21 ➀➁➂➃➇➉ Proposes pairwise self-consistent learning for training CNN to extract these source features and detect deep vacation images.
@@
Finfer hu2022finfer AAAI’22 ➀➂➇zi2020wilddeepfake Based on an autoregressive model, using the facial representation of the current frame to predict the facial representation of future frames.
@@
Huang et alhuang2023implicit CVPR’23 ➀➁➂➉li2019faceshifter A new implicit identity-driven face exchange detection framework is proposed.
@@
HiFi-Net guo2023hierarchical CVPR’23 Converting forgery detection and localization into a hierarchical fine-grained classification problem.
Zhai et alzhai2023towards ICCV’23 dong2013casia hsu2006detecting wen2016coverage Weakly supervised image processing detection is proposed such that only binary image level labels (real or tampered) are required for training.
@@ 3.2 Forgery Detection
In this section, we review current forgery detection techniques based on the type of detection cues, categorizing them into three: Space Domain (Sec. 3.2.1), Time Domain (Sec. 3.2.2), Frequency Domain (Sec. 3.2.3), and Data-Driven (Sec. 3.2.4). We also summarize the detailed information about popular methods in Table 6.
@@ 3.2.1 Space Domain
$\bullet$ Image-level Inconsistency. The generation process of forged images often involves partial alterations rather than global generation, leading to common local differences in non-globally generated forgery methods. Therefore, some methods focus on differences in image spatial details as criteria for determining whether an image is forged, such as color he2019detection , saturation mccloskey2019detecting , artifacts zhao2021multi ; shiohara2022detecting ; cao2022end , gradient variations tan2023learning , etc. Specifically, RECCE cao2022end considers shadow generation from a training perspective, utilizing the learned representations on actual samples to identify image reconstruction differences. LGrad tan2023learning utilizes a pre-trained transformation model, converting images to gradients to visualize general artifacts and subsequently classifying based on these representations. In addition, some works focus on detection based on differences in facial and non-facial regions nirkin2021deepfake , as well as the fine-grained details of image textures chai2020makes ; liu2020global . Recently, Ba et alba2024exposing focuse not only on the discordance in a single image region but also on the detection of fused local representation information from multiple non-overlapping areas.
$\bullet$ Local Noise Inconsistency. Image forgery may involve adding, modifying, or removing content in the image, potentially altering the noise distribution in the image. Detection methods based on noise aim to identify such local or even global differences in the image. Zhou et alzhou2017two propose a dual-stream structure, combining GoogleNet with a triplet network to focus on tampering artifacts and local noise in images. Nguyen et alnguyen2019capsule utilize capsule networks to detect forged pictures and videos in various forgery scenarios. NoiseDF wang2023noise specializes in identifying underlying noise traces left behind in Deepfake videos, introducing an efficient and novel Multi-Head Relative Interaction with depth-wise separable convolutions to enhance detection performance.
@@ 3.2.2 Time Domain
$\bullet$ Abnormal Physiological Information. Forgery videos often overlook the authentic physiological features of humans, failing to achieve overall consistency with authentic individuals. Therefore, some methods focus on assessing the plausibility of the physiological features of the generated faces in videos. Li et alli2018ictu detect blinking and blink frequency in videos as criteria for determining the video’s authenticity. Yang et alyang2019exposing focuses on the inconsistency of head poses in videos, comparing the differences between head poses estimated using all facial landmarks and those estimated using only the landmarks in the central region. Peng et alpeng2024deepfakes focuse on inter-frame gaze angles, obtaining gaze characteristics of each video frame and using a spatio-temporal feature aggregator to combine temporal gaze features, spatial attribute features, and spatial texture features as the basis for detection and classification.
@@
$\bullet$ Inter-Frame Inconsistency. Methods zheng2021exploring ; gu2022delving ; yin2023dynamic ; yang2023masked ; choi2024exploiting ; xu2024towards based on inter-frame inconsistency for forgery detection aim to uncover differences in images between adjacent frames or frames with specific temporal spans. Gu et algu2022delving focuse on inter-frame image inconsistency by densely sampling adjacent frames, while Yin et alyin2023dynamic design a Dynamic Fine-grained Difference Capturing module and a Multi-Scale Spatio-Temporal Aggregation module to cooperatively model spatio-temporal inconsistencies. Yang et alyang2023masked approach forgery detection as a graph classification problem, emphasizing the relationship information between facial regions to capture the relationships among local features across different frames. Choi et alchoi2024exploiting discover that the style variables in each frame of Deepfake work change. Based on this, they developed a style attention module to focus on the inconsistency of the style latent variables between frames.
@@
$\bullet$ Multimodal Inconsistency. The core idea behind multimodal detection algorithms is to make judgments based on the flow of prior information from multiple attributes rather than solely considering the image or audio differences of individual characteristics in each frame. The consideration of audio-visual modal inconsistency has received extensive research in various methods haliassos2022leveraging ; cozzolino2023audio ; feng2023self ; yang2023avoid . POI-Forensics cozzolino2023audio proposes a deep forgery detection method based on audio-visual authentication, utilizing contrastive learning to learn the most distinctive embeddings for each identity in moving facial and audio segments. AVoiD-DF yang2023avoid embeds spatiotemporal information in a spatiotemporal encoder and employs a multimodal joint decoder to fuse multimodal features and learn their inherent relationships. Subsequently, a cross-modal classifier is applied to detect disharmonious operations within and between modalities. Agarwal et alagarwal2021detecting describe a forensic technique for detecting fake faces using static and dynamic auditory ear characteristics. Indeed, multimodal detection methods are currently a hotspot in forgery detection research.